### v1

# GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and A Comprehensive Multimodal Dataset Towards General Medical AI

Abstract
Despite significant advancements in general artificial intelligence, such as GPT-4, their effectiveness in the medical domain (general medical AI, GMAI) remains constrained due to the absence of specialized medical knowledge. To address this challenge, we present GMAI-VL-5.5M, a comprehensive multimodal medical dataset created by converting hundreds of specialized medical datasets into meticulously constructed image-text pairs. This dataset features comprehensive task coverage, diverse modalities, and high-quality image-text data. Building upon this multimodal dataset, we propose GMAI-VL, a general medical vision-language model with a progressively three-stage training strategy. This approach significantly enhances the model’s ability by integrating visual and textual information, thereby improving its ability to process multimodal data and support accurate diagnosis and clinical decision-making. Experimental evaluations demonstrate that GMAI-VL achieves state-of-the-art results across a wide range of multimodal medical tasks, such as visual question answering and medical image diagnosis. Our contributions include the development of the GMAI-VL-5.5M dataset, the introduction of the GMAI-VL model, and the establishment of new benchmarks in multiple medical domains. Code and dataset will be released at https://github.com/uni-medical/GMAI-VL.
1 Introduction
Large-scale Vision-Language Models (LVLMs) have rapidly evolved in recent years, effectively integrating visual perception with language understanding by leveraging large-scale multimodal data, which enables them to capture complex visual and textual patterns and drive significant advancements in image recognition, natural language processing, and multimodal tasks. With the advancement of multimodal integration technologies, the demand for high-precision processing of diverse data types in the medical field has become increasingly critical. The ability to effectively integrate and analyze various data modalities, such as medical images, clinical text, and structured clinical records, is pivotal for achieving accurate and comprehensive diagnostic and treatment decisions.
However, existing LVLMs, such as GPT-4 [1], are limited in medical applications due to their lack of domain-specific knowledge, highlighting the need for specialized solutions that effectively integrate medical expertise. Addressing this challenge requires constructing a comprehensive medical vision-language dataset and developing domain-specific models. For the medical dataset, it should provide high-quality medical knowledge, including the following three aspects:
Comprehensive medical tasks. To enhance the model’s applicability across various medical scenarios, the dataset should cover a wide range of medical contexts, such as disease types, symptoms, and treatments. Comprehensive task coverage can improve the model’s generalization ability and increase its reliability in real-world applications. However, existing models often focus on specific domains [28, 41, 72, 65, 35], limiting their broader applicability. Expanding the dataset’s scope will further enhance the model’s utility in clinical practice.
Rich multimodal representation. A well-rounded medical multimodal dataset should encompass various modalities, including different medical imaging types (such as CT, MRI, and X-rays) and diverse forms of textual data (such as medical records and imaging reports). This would allow models to better integrate multi-source information and improve their analytical capabilities. However, existing methods tend to focus on a single type of medical imagery [33, 70, 50], limiting the model’s adaptability to diverse clinical scenarios. A more diverse multimodal dataset would provide a foundation for developing more comprehensive models, better suited to the complexity of real-world medical environments.
Figure 1: Overview of GMAI-VL and GMAI-VL-5.5M. (a) illustrates the sources, departments, modalities, task types, and instruction formats of the GMAI-VL-5.5M dataset. (b) Architecture of GMAI-VL, integrating a Vision Encoder, Projector, and Large Language Model. (c) Three-stage training process of GMAI-VL, including shallow alignment, deep alignment, and instruction tuning with corresponding data sizes and training components. The flame symbol Refer to caption denotes the training part, while the snowflake symbol Refer to caption indicates frozen part.
High-quality image-text data. High-quality training data is essential for optimizing model performance, especially in medical applications. Ideally, image-text data for medical contexts should consist of a large, well-curated collection of medical images paired with precise textual descriptions. This data would enhance the model’s understanding of key medical concepts, such as diagnosis, treatment, and clinical workflows, ultimately contributing to improved clinical outcomes. While progress has been made with datasets sourced from platforms like PubMed [39, 54, 70], these datasets suffer from inconsistent data quality, imprecise alignment between images and text, and a lack of standardization, which limits their effectiveness.
Based on above observations, we propose an annotation-guided data generation methodology for developing a comprehensive multimodal medical datasets. The methodology begins by collecting large-scale, open-source medical imaging datasets and extracting key annotations, such as modality, task type, labels, and bounding boxes. A vision-language model (e.g., GPT-4o) is then used to transform these datasets, covering tasks like lesion detection, segmentation, and disease diagnosis, into high-quality image-text pairs for training LVLMs. To ensure data quality, extracted image information is incorporated into the prompt design, improving model performance across various clinical tasks. This results in a comprehensive multimodal dataset with 5.5M samples, named GMAI-VL-5.5M, which supports the development of general medical LVLMs. Fig. 1(a) illustrates the sources, departments, modalities, task types, and instruction formats of the constructed dataset.
With the constructed GMAI-VL-5.5M dataset, we develop a general medical vision-language model, GMAI-VL. To enhance its integration of visual and linguistic features and its instruction-following abilities, a three-stage training strategy is proposed in this paper. Specifically, we sequentially implement shallow and deep alignments in the first two stages, gradually building associations between visual (medical images) and language (medical texts) elements from basic features to high-level semantics. Next, we fine-tune the model with cross-modal instructions, improving its understanding of visual-language interactions and instruction-following in complex tasks. With this strategy, GMAI-VL shows strong performance in medical tasks like visual question answering and medical image diagnosis, providing a solid foundation for advancing multimodal models in the medical field.
The contributions of this work are as follows:
  • We build an annotation-guided data generation methodology to construct GMAI-VL-5.5M, a comprehensive vision-language dataset. This dataset offers extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, providing a robust foundation for model training.
  • Leveraging GMAI-VL-5.5M, we design GMAI-VL, a versatile medical vision-language model. Our proposed three-stage training strategy enhances its ability to integrate visual and linguistic features, significantly improving its performance in instruction-following and generalization across a wide range of medical tasks.
  • GMAI-VL outperforms existing models in multimodal question-answering tasks, including PMC-VQA and VQA-RAD. It sets new benchmarks on OmniMedVQA, GMAI-MMBench, and Health & Medicine track of MMMU. Specifically, GMAI-VL achieves an average score of 88.48% on OmniMedVQA, 62.43% on the GMAI-MMBench test set, and 51.3% on the Health & Medicine track of MMMU.
2 Related Work
Large-scale medical vision-language datasets of high quality and multiple modalities are the basis of Large Vision-Language Models (LVLMs) in the medical domain. While natural language and vision datasets are easily accessible online, biomedical datasets often focus on text or images only and many of them are limited to specific tasks or modalities, thus with unsatisfactory generalization ability. Notable datasets like MIMIC-CXR [33] and CheXpert [12] have advanced radiology models but are restricted to single image modality (X-ray), which hinders their use as general-purpose medical LVLMs. To address this, researchers have begun scraping public sources like PubMed and textbooks to construct large-scale vision-language datasets. Examples include datasets proposed in LLaVA-Med [39], Med-Flamingo [54], and PubMedVision [14], with PubMedVision optimizing LLaVA-Med dataset for higher-quality medical data. In addition to scraping, open-source image datasets with annotations can also be converted into image-text pairs for model training. Specifically, image information like modalities and annotations are input into large language models, e.g., GPT series, to generate text paired with the corresponding image. Some popular examples include the datasets constructed in RadFM [70], MedDr [28], MedTrinity-25M [71], ChiMed-VL [48], BiomedGPT [79], Med-Gemini [58], and Med-PaLM [59].
These efforts usually suffer from either limited modalities, data sources, or task coverage. Thus, their dataset quality needs further improvement. To this end, we propose to construct a comprehensive medical vision-language dataset with extensive coverage of medical tasks, diverse multimodal representations, and high-quality image-text pairs, forming a robust foundation for model training.
Medical vision-language models
are usually based on general-purpose Large Vision-Language Models (LVLMs). Most of them adapt LVLMs to specific medical applications using specialized medical datasets. For instance, Med-Flamingo [54] enhances OpenFlamingo-9B using 0.8 million interleaved and 1.6 million paired medical image-text data, highlighting the critical need for multimodal data in medical image analysis and automated report generation tasks. RadFM [70] improves PMC-LLaMA [69] by leveraging 16 million radiology images with text descriptions from diverse sources. Similarly, Med-PaLM [68] adapts PaLM-E [25] to the medical domain with approximately one million medical data samples, achieving state-of-the-art performance in diagnostic support and medical knowledge Q&A. LLaVA-Med [39] utilizes a large-scale biomedical figure-caption dataset extracted from PubMed Central to enhance LLaVA [66, 67] to better understand biomedical images and facilitate open-ended conversational interactions. Med-Gemini [58] leverages long-format question-answering datasets to improve the multimodal and long-contextual capabilities of the baseline Gemini model, enabling superior performance in complex medical Q&A and multimodal reasoning tasks. Additionally, HuatuoGPT-Vision [14] and MedDr [28] build medical multimodal datasets to adapt general-purpose LVLMs like LLaVA and InternVL to various medical modalities, including radiology, pathology, dermatology, and endoscopy.
Previous studies usually focus on constructing medical datasets to adapt general-purpose LVLMs but pay less attention to the adaptation strategies. However, naive training/adaptation strategies may not successfully adapt general-purpose LVLMs to the medical data, due to the large gap between the natural image-text pairs and the medical ones. Moreover, these strategies can hardly align the broad imaging modalities and various types of medical text (e.g., prescriptions, radiology reports, and electronic health records) to obtain generalizable features, thus limiting the models’ performance. Our work thus proposes a novel three-stage training strategy to better integrate the visual and language features to enhance generalization ability.
3 GMAI-VL-5.5M: A Comprehensive Multimodal Dataset
With rapid advancements in medical vision-language models (VLMs), constructing high-quality datasets is crucial for developing general-purpose models. Unlike previous approaches that primarily rely on published literature, our method utilizes specialized medical datasets to create a more robust, high-quality resource. We build GMAI-VL-5.5M, a comprehensive medical vision-language dataset that aggregates data from diverse open-source and proprietary sources. Covering 13 medical imaging modalities and 18 specialties, it addresses a wide range of medical imaging tasks. This dataset enhances the model’s ability to process complex medical information, advancing precision medicine and intelligent diagnostics.
Table 1: Comparison of various medical multimodal datasets, including details on the dataset size, modality type, language, data traceability, and sources of information.
Datasets Data Size Modality Language Traceability Data Source
PathVQA [29] 32.7k Pathology EN × Textbooks
MIMIC-CXR [33] 227k X-Ray EN Hospital
quilt-1M [32] 1M Pathology EN × YouTube & PubMed
MedDr VQA [28] 197k Multimodal EN 13 medical datasets
PMC-OA [43] 1.65M Multimodal EN × PubMed
PMC-VQA [80] 413k Multimodal EN × PubMed
LLaVA-Med VQA [39] 56,702 Multimodal EN × PubMed
ChiMed-VL [48] 1.05M Multimodal CN × PubMed
PMC-CaseReport [70] 438k Multimodal EN × PubMed
PubMedVision [14] 1.29M Multimodal EN&CN × PubMed
GMAI-VL-5.5M (ours) 5.5M Multimodal EN&CN 219 specialized medical imaging datasets
3.1 Data Curation
Data collection.
To construct a comprehensive multimodal medical dataset, we sourced 219 datasets from diverse platforms. Fig. 1(a) highlights key data sources, such as Kaggle, Grand Challenge, and Huggingface, which facilitate extensive data collection. These datasets cover diverse imaging modalities, including fundus, CT, MRI, and ultrasound (US), and span a variety of medical tasks, such as diagnosis, severity assessment, and organ recognition. They also encompass multiple clinical specialties, including pathology, dermatology, ophthalmology, otolaryngology, and oncology, further enhancing their diversity.
Data preprocess.
After data collection, we apply a preprocessing workflow to extract 2D medical images from the videos and 3D medical volumes. For data cleaning, we first extract key annotation information from classification data, including modality, department, and labels for each image, discarding instances with missing or unclear annotations. For segmentation data, we follow the SA-Med2D-20M [75] preprocessing method, filtering out low-quality images and labels, and converting them into detection datasets. The converted segmentation data is then processed together with the original detection dataset, with key annotation information extracted and linked to each image. Finally, the preprocessed data are then standardized and organized into a structured format: <image, modality, label, department, bbox [optional]>.
Data generation.
Subsequently, the data are categorized into two primary types: classification datasets and detection/segmentation datasets. Each category is further refined using specific prompts tailored for large model training. For data generation, large vision-language models (GPT-4o) are employed to produce detailed image descriptions and corresponding instruction-following data through annotation-guided data generation methodology. For classification datasets, detailed descriptions of the entire image are generated, while for detection datasets, the focus is on specific regions enclosed by bounding boxes, providing comprehensive functional analyses of these areas. Notably, the segmentation dataset was transformed into a detection dataset using external bounding boxes, and data generation followed detection dataset protocols. Furthermore, to improve the model’s multilingual capability, we translated a portion of English image-text data into Chinese. Incorporating multilingual data helps to enhance the generalization capabilities of domain-specific multimodal models. The resulted data is utilized for medical Visual Question Answering (VQA) tasks, forming the comprehensive VQA dataset, named GMAI-VL-5.5M. The detailed pipeline for generating prompt-driven data is illustrated in Fig. 2.
Figure 2: The prompt-driven data generation pipeline comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs. Figure with complete prompt and response is provided in Supp. Mat..
As depicted in Fig. 1(a), it contains six kinds of instrcuction-following formats, including image-level captions, region-level captions, free instructions, dialogue, visual perception and text-only tasks. The specific composition of GMAI-VL-5.5M can be found in Supp. Mat. (Table 6). These formats enable VLMs to better understand and process complex visual and textual information in medical contexts. The GMAI-VL-5.5M dataset significantly enhances the model’s cross-modal reasoning ability, enabling it to handle complex multimodal inputs in real clinical scenarios. The richness of the instruction formats allow the model to progress from basic question answering to advanced medical image analysis, ultimately providing strong support for clinical diagnosis and decision-making.
3.2 Data Property
Data statistics.
These datasets encompass diverse medical imaging tasks and modalities, forming a solid foundation for developing and evaluating medical LVLMs. Fig. 4 (in Supp. Mat.) illustrates the distribution of modalities, tasks, clinical departments, and specific medical challenges represented within the collected datasets. This visualization highlights the extensive diversity and coverage of our data collection efforts. After careful standardization and integration, these datasets form the core of our comprehensive medical image-text dataset, GMAI-VL-5.5M, which serves as a crucial resource for advancing precision medicine and intelligent diagnostic systems.
Accuracy and reliability of generated data.
The accuracy and reliability of the generated data are primarily addressed from two key aspects. First, the reliability of the data source is ensured by utilizing high-quality datasets from reputable sources. The data used for generation primarily comes from professional medical challenges (such as Kaggle and GrandChallenge) and publicly available, peer-reviewed datasets. This guarantees the integrity and accuracy of the data at its source. Second, the data generation process is rigorously controlled to ensure accuracy and reliability. Although GPT was used for data generation, prompts were meticulously crafted to include essential annotations, minimizing errors and hallucinations. As illustrated in Fig. 2, the proposed annotation-guided approach incorporates specific annotations (e.g., <image, modality, label, department, bbox [optional]>) to produce precise, high-quality descriptions. Compared to the without-annotation-guided approach, which often results in less detailed outputs, the annotation-guided method delivers more professional, pathology-specific descriptions, significantly enhancing the reliability of the generated data.
Compared with other medical multimodal datasets.
The GMAI-VL-5.5M dataset, as highlighted in Table 1, stands out due to its unmatched scale, encompassing over 5.5 million samples from more than 219 specialized medical imaging datasets. Unlike other datasets listed, GMAI-VL-5.5M supports a wider variety of modalities and languages, making it a truly global resource that caters to diverse clinical needs. Additionally, GMAI-VL-5.5M emphasizes traceability of its data, ensuring a high standard of clinical relevance and reliability. This comprehensive and diverse dataset is critical for pushing the boundaries of medical multimodal research, enabling more effective training of LVLMs that can generalize across multiple medical tasks and scenarios, thereby driving innovations in precision medicine and intelligent diagnostics.
4 GMAI-VL: A General Medical Vision-Language Model
4.1 Architecture
The GMAI-VL model is a vision-language model built upon the LLaVA architecture [45, 39], incorporating three key components: a large language model (LLM), a vision encoder, and a projector (MLP), as illustrated in Fig. 1(b). These components are designed to work together seamlessly, enabling the model to deliver exceptional performance in medical applications.
We adopt InternLM2.5-7B [63] as our language processing module, which offers outstanding reasoning capabilities. With a context window up to one million tokens, it can handle complex medical tasks and generate coherent, accurate responses. Its support for advanced instruction-following makes it particularly effective in addressing intricate medical queries, thereby enhancing the model’s ability to understand and respond to a wide range of instructions.
For vision processing, GMAI-VL employs a CLIP-based vision encoder [55], which transforms visual inputs into high-dimensional feature representations. CLIP’s strong performance in aligning image and text representations ensures that medical image features are accurately extracted and effectively integrated with linguistic information, significantly enhancing the model’s ability to handle multimodal medical data.
The MLP, as a projector, serves as a bridge between the vision encoder and the LLM, optimizing high-dimensional outputs and further enhancing feature representation. The seamless integration of these components enables GMAI-VL to excel in processing and understanding complex medical multimodal data.
4.2 Optimization Strategy
As illustrated in Fig. 1(c), the training process of the GMAI-VL model is divided into three stages: shadow alignment, deep alignment, instruction tuning, respectively. The detailed hyper-parameter settings can be found in Supp. Mat. (Table 8). To enhance the training of GMAI-VL, we supplement our GMAI-VL-5.5M dataset with additional medical datasets, ultimately forming an 11.7M dataset for training. This supplemented data increases the diversity of training data, exposing the model to a wider range of medical scenarios and visual-language patterns, enhancing its generalization to complex clinical tasks and ensuring robustness in real-world applications. Fig. 5 (in Supp. Mat.) provides a complete distribution of the utilized datasets during the training stage. The detailed data proportions for each training stage are detailed in Table 7 (in Supp. Mat.).
Stage I: Shadow alignment.
In the shallow alignment phase, we utilize a large-scale medical image-text dataset comprising approximately 11.7 million image-text pairs, sourced from a combination of publicly accessible datasets and proprietary in-house data. To achieve shallow alignment, we freeze both the large language model and the vision encoder, optimizing only the projector. With this optimization stage, the model establishes an initial alignment between medical images and their corresponding textual descriptions. All input images are resized to $336\times 336$ pixels, and the training objective is to minimize the cross-entropy loss of the text tokens.
Stage II: Deep alignment.
Most vision encoders in multimodal models are pre-trained on natural images, so it is essential to address the domain gap between medical and natural images during the deep alignment stage. To achieve this, we fine-tune both the vision-language projector and the vision encoder, enhancing the alignment between the visual features of medical images and the language model’s feature space.
Stage III: Instruction tuning.
At this stage, we fine-tune our GMAI-VL model (including the vision encoder, the language model, and the projector components) by instruction tuning to enhance its instruction-following and dialogue capabilities. The multimodal instruction tuning data is primarily derived from the training data in previous stages, with a filtering process to select high-quality and more suitable data for fine-tuning. Specifically, we manually reviewed a subset of each dataset and labeled it as either “high quality” or “low quality.” High-quality datasets were used for tuning, while most low-quality ones were excluded, except for a small portion retained for diversity. This approach allowed efficient filtering without reviewing every sample. Additionally, we incorporate medical text dialogue data to ensure the model’s versatility in handling various dialogue scenarios. Thus, our instruction tuning data comprises approximately ten million samples.
5 Experimental Results
To evaluate our model, we employed several established multimodal medical benchmarks, each targeting specific aspects of medical image understanding and question answering. Below is a brief overview of the benchmarks used in our experiments:
  • Traditional medical VQA benchmarks: Traditional multimodal medical question-answering benchmarks, such as VQA-RAD [38], SLAKE [44], and PMCVQA [80], span various imaging modalities tasks, primarily assessing the model’s ability to extract information from medical images and answer clinical questions. They evaluate the model’s performance in understanding medical imaging and integrating multimodal information.
  • OmniMedVQA: OmniMedVQA [31] provides a rich dataset of paired medical images and text, designed to evaluate the model’s ability to recognize and understand fundamental medical imaging concepts, focusing on cross-modal reasoning and information integration.
  • GMAI-MMBench: GMAI-MMBench [16] focuses on assessing the model’s ability to identify fine-grained objects in complex clinical scenarios, challenging its capacity to handle long-context tasks and accurately recognize and reason over detailed medical features.
  • MMMU Health & Medicine track: The Health & Medicine track of the MMMU [78] benchmark spans a wide range of medical fields, derived from university exams, quizzes, and textbooks. It evaluates the model’s reasoning ability in complex medical scenarios and the specialized knowledge in health and medicine.
5.1 Medical VQA Benchmarks
Table 2: Results on Traditional Medical VQA Benchmarks. The highest performance in each column is highlighted in red, and the second-highest performance is highlighted in blue.
Model VQA-RAD SLAKE PMC-VQA Avg.
Med-Flamingo [54] 45.4 43.5 23.3 37.4
RadFM [70] 50.6 34.6 25.9 37.0
LLAVA-Med-7B [39] 51.4 48.6 24.7 41.6
Qwen-VL-Chat [6] 47.0 56.0 36.6 46.5
Yi-VL-34B [77] 53.0 58.9 39.5 50.5
LLAVA-v1.6-7B [46] 52.6 57.9 35.5 48.7
LLAVA-v1.6-13B [46] 55.8 58.9 36.6 50.8
LLAVA-v1.6-34B [46] 58.6 67.3 44.4 56.8
HuatuoGPT-Vision-7B [14] 63.8 74.5 52.7 63.7
GMAI-VL(w/o our data) 62.3 66.3 39.0 55.9
GMAI-VL(ours) 66.3 72.9 54.3 64.5
The performance of various VLMs on popular medical VQA benchmark datasets is summarized in Table 2, including VQA-RAD[38], SLAKE[44], and PMC-VQA[80]. Our model, GMAI-VL, demonstrates a strong performance, achieving the highest score on the VQA-RAD[38] dataset with 66.3%, outperforming other models such as HuatuoGPT-Vision-7B. This result highlights GMAI-VL’s superior capability in handling radiological image question-answering tasks. For the PMC-VQA[80] dataset, GMAI-VL achieves 54.3%, and 72.9% on SLAKE[44], demonstrating its capability in handling medical VQA tasks across diverse modalities.
GMAI-VL demonstrates competitive performance across multiple benchmarks, showcasing its versatility in medical image understanding and question-answering.
5.2 OmniMedVQA
Table 3: Comparison of performance between representative LVLMs and GMAI-VL on OmniMedVQA across five different question type. The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue. Abbreviations: MR = Modality Recognition, AI = Anatomy Identification, DD = Disease Diagnosis, LG = Lesion Grading, OBA = Other Biological Attributes.
Model MR AI DD LG OBA Overall
Random Guess 25.00 25.84 28.41 25.40 37.49 28.28
Open-Source LVLMs
MiniGPT-4 [81] 36.98 32.68 24.19 20.45 26.14 27.59
LLaVA [45] 52.30 35.27 11.80 9.77 24.70 22.86
LLaMA_Adapter_v2 [27] 58.45 38.18 29.12 23.73 30.97 35.08
InstructBLIP [20] 72.35 39.90 32.01 43.80 47.91 41.14
BLIP-2 [40] 57.48 49.83 46.21 30.52 73.52 50.77
Qwen-VL-Chat [6] 33.69 10.95 16.27 6.71 41.68 20.29
mPLUG-Owl2 [76] 78.01 48.52 39.68 20.56 59.36 48.44
LLaVa-NeXT [46] 68.23 46.74 41.21 18.43 39.57 45.57
DeepSeek-VL [49] 74.01 51.94 45.46 21.06 29.04 48.76
Yi-VL [77] 59.56 44.81 48.97 32.93 24.63 47.28
InternVL2-40B [18] 96.76 64.25 76.28 76.50 76.27 78.70
Medical Special Model
MedVInT-TE [80] 62.62 41.03 40.57 12.17 45.17 43.83
LLaVA-Med [39] 48.41 27.96 23.72 16.10 21.94 27.82
Med-Flamingo [54] 26.74 25.10 23.80 28.04 16.26 23.82
RadFM [70] 27.45 21.65 23.75 16.94 20.05 23.48
MedDr [28] 91.37 51.62 65.56 73.18 74.52 68.27
HuatuoGPT-Vision-34B [14] 95.06 75.67 66.51 72.83 74.92 73.23
Our Model
GMAI-VL(w/o our data) 96.40 80.97 79.14 70.29 75.66 79.96
GMAI-VL(ours) 98.64 92.95 88.7 87.21 82.95 88.48
Table 3 summarizes the performance of various large vision-language models (LVLMs), including our proposed GMAI-VL, across five question types: Modality Recognition, Anatomy Identification, Disease Diagnosis, Lesion Grading, and Other Biological Attributes. GMAI-VL demonstrates outstanding accuracy across multiple tasks, achieving 98.64% in Modality Recognition, 92.95% in Anatomy Identification, and 88.71% in Disease Diagnosis. It outperforms both open-source LVLMs and medical-specific models, underscoring its capability to accurately identify anatomical structures and diagnose diseases from visual data. In Lesion Grading, GMAI-VL attained the highest score of 87.21%, and it also delivers a strong performance of 82.95% in Other Biological Attributes, showcasing its versatility across diverse biological contexts. With an average accuracy of 88.48%, the highest among all evaluated models, GMAI-VL excels not only in general medical question-answer tasks but also in complex reasoning requiring domain-specific knowledge, surpassing models like HuatuoGPT-Vision-34B and InternVL2-40B.
These results verify our GMAI-VL is a leading model in multimodal medical image understanding, setting a new benchmark for medical VQA tasks. Its consistent top performance across question types highlights its potential for broader applications in medical question answering.
5.3 GMAI-MMBench
Table 4: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks. The full names of the evaluated tasks can be found in Table 5 in literature [16]. The best model in each category is highlighted in red, while the second-best model is indicated in blue.
Model Name Overall (val) Overall (test) AR BVR B CR C DD IQG MR M NT OR-A OR-HN OR-P OR-T SG SAR SIR SWR
Random Guess
Random 25.70 25.94 38.20 22.73 22.92 22.72 24.06 26.66 27.13 27.00 20.00 24.75 21.37 22.93 22.33 21.18 32.43 24.23 21.39 23.71
Open-Source LVLMs
Flamingo v2 [4] 25.58 26.34 37.74 21.50 20.62 22.00 22.41 27.29 25.91 27.45 18.00 28.79 25.16 22.13 22.00 22.00 34.61 22.88 20.44 27.43
VisualGLM-6B [22] 29.58 30.45 40.16 33.92 24.92 25.22 24.21 32.99 29.96 29.53 21.20 37.88 30.32 24.80 13.33 29.88 33.11 19.62 19.16 37.43
InstructBLIP-7B  [20] 31.80 30.95 42.12 26.92 24.92 28.09 21.65 34.58 31.58 29.23 22.40 30.30 28.95 27.47 23.00 24.82 32.88 19.81 21.64 26.57
Qwen-VL  [6] 34.80 36.05 37.05 37.24 35.85 28.98 24.81 43.60 24.70 30.12 19.20 44.44 29.68 31.87 25.00 31.18 30.26 21.54 20.10 26.86
Yi-VL-6B  [77] 34.82 34.31 41.66 39.16 26.62 30.23 31.88 38.01 26.72 24.93 25.20 37.37 29.58 31.20 32.33 30.59 36.71 24.81 23.18 31.43
ShareGPT4V-7B  [15] 36.71 36.70 43.96 37.59 21.54 37.57 18.80 43.26 32.39 27.30 22.80 43.43 29.47 37.33 22.00 31.76 34.98 24.42 25.06 30.00
LLAVA-V1.5-7B [45] 38.23 37.96 45.45 34.27 30.92 41.32 21.65 44.68 34.01 27.74 23.60 43.43 28.00 42.13 29.00 35.06 33.41 22.12 23.61 29.14
XComposer2  [24] 38.68 39.20 41.89 37.59 33.69 40.79 22.26 45.87 36.44 32.94 27.20 58.59 26.11 36.40 43.67 37.29 32.06 23.46 27.80 32.86
LLAVA-InternLM-7b [19] 38.71 39.11 36.36 36.54 32.62 38.10 30.68 46.53 34.82 28.19 25.20 48.99 28.11 40.53 33.33 36.00 34.08 26.73 24.12 29.71
InternVL-Chat-V1.5 [18] 38.86 39.73 43.84 44.58 34.00 33.99 31.28 45.59 33.20 38.28 32.40 42.42 31.89 42.80 27.00 36.82 34.76 23.27 24.72 32.57
InternVL-Chat-V1.2  [17] 39.52 40.01 41.66 44.06 27.38 38.46 34.29 46.99 33.60 34.42 21.20 47.98 30.63 42.80 27.67 35.88 35.59 23.85 24.98 28.00
LLAVA-InternLM2-7b [19] 40.07 40.45 39.82 37.94 30.62 35.24 29.77 48.97 34.01 25.96 20.80 53.03 30.95 42.67 32.00 39.88 32.43 21.73 24.38 38.00
DeepSeek-VL-7B  [49] 41.73 43.43 38.43 47.03 42.31 37.03 26.47 51.11 33.20 31.16 26.00 44.95 36.00 58.13 36.33 47.29 34.91 18.08 25.49 39.43
MiniCPM-V2 [73] 41.79 42.54 40.74 43.01 36.46 37.57 27.82 51.08 28.74 29.08 26.80 47.47 37.05 46.40 25.33 46.59 35.89 22.31 23.44 31.71
Proprietary LVLMs
Claude3-Opus [2] 32.37 32.44 1.61 39.51 34.31 31.66 12.63 39.26 28.74 30.86 22.40 37.37 25.79 41.07 29.33 33.18 31.31 21.35 23.87 4.00
Qwen-VL-Max  [6] 41.34 42.16 32.68 44.58 31.38 40.79 10.68 50.53 32.79 44.36 29.20 51.52 41.37 58.00 30.67 41.65 26.95 25.00 24.64 39.14
GPT-4V  [1] 42.50 44.08 29.92 48.95 44.00 37.39 12.93 52.88 32.79 44.21 32.80 63.64 39.89 54.13 37.00 50.59 27.55 23.08 25.75 37.43
Gemini 1.0  [62] 44.38 44.93 42.12 45.10 46.46 37.57 20.45 53.29 35.22 36.94 25.20 51.01 34.74 59.60 34.00 50.00 36.64 23.65 23.87 35.43
Gemini 1.5  [56] 47.42 48.36 43.50 56.12 51.23 47.58 2.26 55.33 38.87 48.07 30.00 76.26 51.05 75.87 46.33 62.24 20.57 27.69 30.54 40.57
GPT-4o  [1] 53.53 53.96 38.32 61.01 57.08 49.02 46.62 61.45 46.56 56.38 34.00 75.25 53.79 69.47 48.67 65.88 33.93 22.88 29.51 39.43
Medical Special Model
Med-Flamingo  [54] 12.74 11.64 6.67 10.14 9.23 11.27 6.62 13.43 12.15 6.38 8.00 18.18 9.26 18.27 11.00 11.53 12.16 5.19 8.47 11.43
LLaVA-Med  [39] 20.54 19.60 24.51 17.83 17.08 19.86 15.04 19.81 20.24 21.51 13.20 15.15 20.42 23.73 17.67 19.65 21.70 19.81 14.11 20.86
Qilin-Med-VL-Chat  [48] 22.34 22.06 29.57 19.41 16.46 23.79 15.79 24.19 21.86 16.62 7.20 13.64 24.00 14.67 12.67 15.53 26.13 24.42 17.37 25.71
RadFM  [70] 22.95 22.93 27.16 20.63 13.23 19.14 20.45 24.51 23.48 22.85 15.60 16.16 14.32 24.93 17.33 21.53 29.73 17.12 19.59 31.14
MedDr [28] 41.95 43.69 41.20 50.70 37.85 29.87 28.27 52.53 36.03 31.45 29.60 47.47 33.37 51.33 32.67 44.47 35.14 25.19 25.58 32.29
Our Model
GMAI-VL(w/o our data) 54.99 56.23 51.26 61.05 53.79 44.39 44.51 62.60 40.80 57.42 35.20 79.50 61.31 77.81 53.60 69.29 35.39 35.77 29.71 44.86
GMAI-VL(ours) 61.74 62.43 75.26 59.66 67.24 56.86 54.29 67.14 42.80 79.97 41.60 75.00 60.45 75.48 53.33 58.12 42.09 72.31 37.40 59.14
The GMAI-MMBench benchmark is a comprehensive medical multimodal benchmark designed to evaluate models on a range of clinical visual question-answering (VQA) tasks. Table 4 presents the results of various LVLMs, including open-source LVLMs and commercial models, evaluated on the val and test sets across multiple clinical tasks.
Notably, GMAI-VL excels in specific tasks such as abnormality recognition (73.78%), biological variation recognition (63.06%), and clinical disease diagnosis (66.67%). These results demonstrates the model’s strong ability in understanding and interpreting complex clinical images. In comparison to other models, GMAI-VL consistently achieves either the best or second-best performance across most tasks. For instance, it ranks first in 16 out of 20 categories, including key tasks such as AR (Attribute Recognition) and DD (Disease Diagnosis), where it achieved scores of 75.26% and 67.14%, respectively, suggesting GMAI-VL’s strength in understanding medical scenarios.
Overall, GMAI-VL establishes a new benchmark in various clinical VQA tasks, demonstrating its potential as a reliable and versatile tool in medical multimodal applications.
5.4 MMMU Health & Medicine track
The MMMU benchmark, a widely recognized standard for evaluating multimodal models, was used to assess the performance of our proposed GMAI-VL model on the Health & Medicine track. The experimental results in Table 5 show the model’s performance across five key categories: Basic Medical Science (BMS), Clinical Medicine (CM), Diagnostics and Laboratory Medicine (DLM), Pharmacy (P), and Public Health (PH). GMAI-VL performs strongly across multiple categories, achieving top scores in DLM (43.3%), P (50.0%), and PH (53.3%), surpassing competitive models like LLaVA-v1.6 and HuatuoGPT-Vision-7B. These results highlight the model’s proficiency in handling complex tasks requiring diagnostic reasoning, pharmaceutical knowledge, and public health expertise. In BMS, GMAI-VL scores 50.0%, achieve the best performance, demonstrating the model’s the capacity of understanding medical knowledge. In CM, the model achieves 60.0%, remaining competitive with other leading models. These results underscore the model’s ability in processing both clinical and foundational medical information effectively. Overall, GMAI-VL achieves an average score of 51.3% across the Health & Medicine track, which is a top performance among other models, verifying its versatility in specialized medical domains.
Table 5: Performance on the val set for the MMMU Health & Medicine track. This track is divided into five categories: BMS (Basic Medical Science), CM (Clinical Medicine), DLM (Diagnostics and Laboratory Medicine), P (Pharmacy), and PH (Public Health). The best performance in each column is highlighted in red, and the second-best performance is highlighted in blue.
Model BMS CM DLM P PH MMMU Health & Medicine
Med-Flamingo [54] 33.6 30.2 23.3 29.3 25.8 28.4
RadFM [70] 31.6 28.6 26.7 26.2 26.8 27.9
LLaVA-Med-7B [39] 33.8 32.3 26.7 40.7 43.3 38.6
Qwen-VL-Chat [6] 32.7 20.6 19.3 29.6 33.3 31.7
Yi-VL-34B [77] 48.1 55.6 36.7 35.4 31.3 48.2
LLaVA-v1.6-7B [45] 46.4 43.4 30.0 29.6 26.7 33.1
LLaVA-v1.6-13B [45] 53.6 46.7 33.3 22.2 40.0 39.3
HuatouGPT-Vision-7B [14] 50.0 63.3 36.7 48.1 53.3 50.3
GMAI-VL(w/o our data) 43.3 56.7 43.3 46.7 40.0 46.0
GMAI-VL(ours) 50.0 60.0 43.3 50.0 53.3 51.3
5.5 Ablation Study on GMAI-VL-5.5M Dataset
We further analyzed the effectiveness of the proposed GMAI-VL-5.5M dataset. By removing GMAI-VL-5.5M from the training set, we observed a significant performance drop, as shown in Table 2, Table 3, Table 4 and Table 5. In the VQA-RAD, SLAKE, and PMC-VQA benchmarks, removing GMAI-VL resulted in an average performance improvement of 8.6%. For the OmniMedVQA, GMAI-MMBench, and MMMU Health & Medicine track, the improvements were 8.52%, 6.75%, and 5.3%, respectively. These results demonstrate that the GMAI-VL-5.5M dataset provides highly accurate and reliable medical knowledge, particularly in recognizing and understanding multimodal medical data, which significantly boosts model performance. This underscores the dataset’s diversity, comprehensiveness, and its ability to effectively complement other datasets in complex medical tasks.
6 Conclusion
In this study, we develop GMAI-VL, a large vision-language model, along with GMAI-VL-5.5M, a comprehensive multimodal medical dataset aimed at advancing general medical AI (GMAI). GMAI-VL-5.5M, which converts hundreds of medical image analysis datasets into high-quality image-text pairs through annotation-guided data generation, enables GMAI-VL to effectively address a wide range of clinical tasks. Experimental results show that GMAI-VL-5.5M significantly enhances GMAI-VL’s performance on diverse clinical tasks, achieving state-of-the-art results across several key benchmarks.
Supplementary Material
7 Expanded Data Generation Example
In this section, we present the full version of Fig. 2, which provides a detailed example of the data generation process, as shown in Fig.3. This expanded illustration includes the complete prompt and the corresponding detailed responses, enabling a clear comparison between data generation methodologies with and without annotation guidance. The detailed comparison allows for a more intuitive understanding of the distinctions between both methods.
Figure 3: The full version of Fig. 2 in the main text illustrates the complete of data generation pipeline, comparing without-annotation-guided and annotation-guided methods. The annotation-guided approach integrates specific annotation information (e.g., <image, modality, label, department, bbox [optional]>) to generate high-quality, accurate descriptions, while the without-annotation-guided approach often results in lower-quality outputs.
8 Details of GMAI-VL-5.5M
In Table. 6, we provide sub-datasets information of the multimodal dataset GMAI-VL-5.5M we have constructed. Based on the different data forms introduced in the paper, we have categorized the data into five distinct sub-datasets. These include GMAI-MM-Caption-1.7M, GMAI-MM-Instruct-0.9M, GMAI-MM-Percept-1.3M, GMAI-Text-Single-1M, and GMAI-Text-0.6M. Each sub-dataset corresponds to specific components: image caption data, free instruction data, visual perception data, text-only, and conversation. Furthermore, Fig. 4 presents a comprehensive distribution of the data across different modalities, original tasks, departments, and clinical task within this multimodal dataset, highlighting its extensive richness and diversity.
Table 6: Sub-Dataset Details for GMAI-VL-5.5M
Dataset Sub-Dataset Name Description Size
GMAI-VL-5.5M GMAI-MM-Caption-1.7M A curated set of detailed medical image captions. 1.7M
GMAI-MM-Instrunct-0.9M A diverse set of instructions for medical image analysis. 0.9M
GMAI-MM-Percept-1.3M A dataset of labels for medical image classification and segmentation. 1.3M
GMAI-Text-Single-1M A set of single-round medical dialogues on patient queries 1.0M
GMAI-Text-Multi-0.6M A dataset of multi-turn medical conversations on various topics. 0.6M
9 Training Data Overview
In this section, we provide a comprehensive overview of all datasets utilized for training the GMAI-VL model. The details include the dataset names, their corresponding categories, the amount of data used for training, and the proportion of training data allocated to each dataset during the three phases of model training. Fig. 5 visualizes the distribution of all our training data, showing the proportion of each category and subcategory.
Table. 7 summarizes the datasets employed, highlighting their respective categories and sizes. It is important to note that for certain datasets, we performed data cleaning and bilingual translation. As a result, the dataset sizes reported here may differ from the official numbers.
(a) (a) Modality distribution
(b) (b) Original task distribution
(c) (c) Department distribution
(d) (d) Clinical task distribution
Figure 4: Distribution of GMAI-VL-5.5M across tasks, modalities, departments, and clinical tasks. (a) Original Task Distribution: The dataset includes 2D Classification (50.4%), 3D Segmentation (30.3%), 2D Segmentation (12.7%), and 2D Detection (6.6%). (b) Modality Distribution: In addition to CT (26.8%) and MR (24.7%), X-ray (12.6%), Pathology (11.2%), and less common modalities like Dermoscopy (3.5%), Microscopy (2.4%), and PET (0.2%) are represented. (c) Department Distribution: While Orthopedic Surgery (12.9%) and General Surgery (10.3%) are the top contributors, departments like Endocrinology (1.3%), Infectious Diseases (0.8%), and Urology (0.7%) also provide data. (d) Clinical Task Distribution: Besides Disease Diagnosis (40.4%) and Organ Recognition (16.0%), tasks such as Muscle Recognition (3.3%), Nervous Tissue Recognition (1.5%), and Microorganism Recognition (1.2%) are included. Note: The distribution statistics shown here pertain only to the multimodal components of GMAI-VL-5.5M.
Figure 5: Distribution of all our training data. The inner ring represents major categories, each depicted in a different color. The outer ring corresponds to the subcategories within each major category. The size of each segment is proportional to the amount of data, as indicated in the legend, where the data volume for each subcategory is also provided.
Table 7: List of datasets used in our model. We employ a large collection of image-text data and instruction data for training stage.
Dataset Category Dataset Name Size ratio in stage 1&2 ratio in stage 3
General Captioning ALLaVA[13] 468k 100.0% 50.0%
ShareGPT4V[15] 102k
Medical Captioning GMAI-MM-Caption-1.7M 1.7M 100.0% 100.0%
PubMedVision[14] 1.3M
MedICaT[60] 173k 100.0% 5.0%
MPx-Single[70] 31k
PMC-OA[43] 1.3M
QUILT-1M[32] 643k
Retina Image Bank[3] 22k
Report Generation CheXpertPlus[12] 223k 100.0% 30.0%
MIMIC-CXR[33] 486k
OpenI[21] 7k
General Instruction GeoQA+[11] 72k 100.0% 75.0%
AI2D[36] 12k
SynthDoG[37] 29k
ChartQA[51] 18k
MMChemExam[42] 219k
LLaVA-Instruct-150K[45] 157k
DVQA[34] 200k
DocVQA[52] 10k
Medical Instruction GMAI-MM-Percept-1.3M 1.3M 100.0% 100.0%
GMAI-MM-Instruct-0.9M 0.9M
PubMedVision[14] 1.28M
LLaVA-Med-60k[39] 56k
PMC-Inline[70] 288k 100.0% 10.0%
VQA-Med-2019[8] 3.2k
Medical-Diff-VQA[30] 260k
PathVQA[29] 2.6k
PMC-CaseReport[70] 109k
PMC-VQA[80] 251k
ROCOV2[57] 60k
SLAKE[44] 0.6k
VQA-RAD[38] 0.3k
General Text blossom_orca[5] 20k 0.0% 100.0%
COIG-CQIA[7] 14.8k
Cosmopedia-100k[9] 33k
ShareGPT4V[15] 26k
Orca-Math[53] 379k
Leetcode[10] 1.7k
LogiQA[47] 12.7k
Lima[26] 83k
Open Hermes 2.5[64] 200k
Firefly[74] 189k
UltraChat[23] 189k
Alpaca-Instruct-52K[61] 49k
Medical Text GMAI-Text-Single-1M 1.0M 0.0% 100.0%
GMAI-Text-Multi-0.6M 649k
Overall - 15.7M - -
Figure 6: Diagram of the three-stage training process.
10 Model Training Settings
Table 8: Training settings of GMAI-VL’s stage I, stage II, and stage III.
Settings Stage I Stage II Stage III
freeze LLM True True False
freeze MLP False False False
freeze Vision Encoder True False False
packing type soft packing soft packing soft packing
learning rate 1e-3 1e-4 1e-5
learning rate schedule cosine decay cosine decay cosine decay
optimizer AdamW AdamW AdamW
optimizer hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$
input size 336x336 336x336 336x336
total batch size 32x8x2 32x4x4 32x4x4
drop rate 0.0 0.0 0.0
numerical precision DeepSpeed bf16 DeepSpeed bf16 DeepSpeed bf16
GPUs for training 32xA100 (80G) 32xA100 (80G) 32xA100 (80G)
The table. 8 presents the training settings for GMAI-VL across three stages, detailing key hyperparameters. The train diagram as illustrated in Fig. 6.
  • Stage I (Shallow alignment). During this stage, both the large language model (LLM) and the vision encoder are frozen, while the MLP remains trainable. The learning rate is set to $1e^{-3}$ with a cosine decay schedule, using AdamW as the optimizer. Input size is $336\times 336$, and the total batch size is $32\times 8\times 2$. Here, 32 refers to the number of GPUs used, 8 represents the micro-batch size per GPU, and 2 is the number of gradient accumulation steps. Notably, we use a soft packing technique, which allows each sample to actually contain multiple sequences—estimated at over 2 per sample.
  • Stage II (Deep alignment). The LLM remains frozen, while the vision encoder and MLP are unfrozen and trainable, allowing for deeper feature alignment. The learning rate is lowered to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
  • Stage III (Instruction tuning). In this stage, all components of the model are trainable, allowing for comprehensive fine-tuning. The learning rate is further reduced to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
Each stage utilizes DeepSpeed for mixed-precision training (bf16) on 32 A100 (80GB) GPUs. To accelerate the training rate of the model, we utilize the soft packing method in xtuner***https://github.com/InternLM/xtuner, which packs multiple sequences into each sample to optimize memory usage and efficiency..
Figure 7: Examples of our model in various medical tasks, including image description, disease diagnosis, diagnostic report generation, free question answering, etc. The results demonstrate the model’s versatility across different imaging modalities and tasks, with support for both Chinese and English.
11 Results
Fig. 7 illustrates several examples of our model on various tasks including image description, disease diagnosis, diagnostic report generation, free question answering, etc. These examples cover a diverse range of medical imaging modalities and tasks, highlighting the model’s versatility and general capabilities within medical domain. Furthermore, GMAI-VL supports both Chinese and English, thereby facilitating bilingual functionality and expanding its applicability in diverse clinical settings.